What change in potential energy does a 12 ?C charge experience if it moves between point A and point B which have a potential difference of 70.0 V? Express your answer in joules and then convert it to electronvolts (eV).
Answer #1

charge ,

q = 12 ?C

potential difference , V = 70 V

change in potential energy = q * V

change in potential energy = 12 *10^-6 * 70

change in potential energy = 0.00084 J

change in potential energy = (.00084)/(1.602 *10^-19)

change in potential energy = 5.24 *10^15 eV

the change in potential energy is 0.00084 J and 5.24 *1015 eV